OpenStudy (anonymous):

8x+4y=0 7x+8y=4 srry stalker :(

OpenStudy (mertsj):

Multiply the top equation by -2 and add the result to the bottom equation.

OpenStudy (anonymous):

\[4y=-8x\] \[substituting this into the second equation\] \[7x+2(4y)=4\] \[7x+2(-8x)=4\] \[7x-16x=4\] \[x=-4/9\]

OpenStudy (anonymous):

now substitute the value of x into 4y=-8x

OpenStudy (anonymous):

-1,2

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

y=8/9

OpenStudy (anonymous):

i got that but i got it rong the correct answer says its -1,2